Question reads:
"Complete the square on the hyperbola: x^2-4y^2+2x+8y-19=0. Finally, draw a quick sketch. A sketch from a graphics calculator is acceptable."
Don't know to complete the square or what to punch in on "y=" on my graphing calculator! Thanks a bunch!
Zach

Answer by stanbon(75887)

"Complete the square on the hyperbola: x^2-4y^2+2x+8y-19=0. Finally, draw a quick sketch. A sketch from a graphics calculator is acceptable."
Don't know to complete the square or what to punch in on "y=" on my graphing calculator!
----------------------
x^2-4y^2+2x+8y-19=0
Complete the square:
x^2+2x+1 -4(y^2-2y+1) = 19 + 1 -4
----
(x+1)^2 -4(y-1)^2 = 16
----------------------------
Divide by 16 to get:
(x+1)^2/16 - (y-1)^2/4 = 1
----
The graph is a hyperbola opening to the right and left
with center at (-1,1)
----
Yes, it's a challenge to put that in a graphing calculator.
Check with your friends and/or you teacher.
-------
Cheers,
Stan H.